# Deep Stochastic Kinematic Models for Probabilistic Motion Forecasting in Traffic

### v2

@@ Abstract
In trajectory forecasting tasks for traffic, future output trajectories can be computed by advancing the ego vehicle’s state with predicted actions according to a kinematics model. By unrolling predicted trajectories via time integration and models of kinematic dynamics, predicted trajectories should not only be kinematically feasible but also relate uncertainty from one timestep to the next. While current works in probabilistic prediction do incorporate kinematic priors for mean trajectory prediction, variance is often left as a learnable parameter, despite uncertainty in one time step being inextricably tied to uncertainty in the previous time step. In this paper, we show simple and differentiable analytical approximations describing the relationship between variance at one timestep and that at the next with the kinematic bicycle model. These approximations can be easily incorporated with negligible additional overhead into any existing trajectory forecasting framework utilizing probabilistic predictions, whether it is autoregressive or one-shot prediction. In our results, we find that encoding the relationship between variance across timesteps works especially well in unoptimal settings, such as with small or noisy datasets. We observe up to a 50% performance boost in partial dataset settings and up to an 8% performance boost in large-scale learning compared to previous kinematic prediction methods on SOTA trajectory forecasting architectures out-of-the-box, with no fine-tuning. In this paper, we show four analytical formulations of probabilistic kinematic priors which can be used for any Gaussian Mixture Model (GMM)-based deep learning models, quantify the error bound on linear approximations applied during trajectory unrolling, and show results to evaluate each formulation in trajectory forecasting.
@@ I Introduction
Figure 1: Motivating example for probabilistic kinematic priors. In real-world traffic, vehicles have a constrained range of behaviors. For instance, it is not possible for a vehicle to move side to side directly, and driving in reverse is highly unlikely. Without kinematic priors, neural networks may search the space of all trajectories, possible or impossible. Without accounting for analytical variances in trajectories as in previous work, the range of possible future trajectories may also be unrealistic. A vehicle’s immediate trajectory milliseconds into the future may have much less uncertainty than a few seconds into the future.
@@
    A simple and effective method for incorporating analytically-derived kinematic priors into probabilistic models for trajectory forecasting (Section IV), which boosts performance and generalization (Section V) and requires negligible additional overhead computation;
@@ II-B Physics-based Priors for Learning
In our method, we pursue the latter. Since we are not modeling complex systems such as cloth or fluid, simulation of traffic agent states require only a simple, fast, and differentiable update. In addition, since the kinematic models do not describe interactions between agents, the complexity of the necessary model is greatly reduced. Many current SOTA trajectory prediction models are designed intentionally to study the relationships between agents via self-attention; thus, in our method, we leave the complex tasks to the network, and model the simple tasks (e.g., how a vehicle moves forward) with equations.
@@ IV Methodology
In this section, we discuss in detail four different kinematic formulations as priors for current trajectory forecasting methods.
@@ IV-B1 Formulation 1: Velocity Components
Figure 4: Average Displacement Error (ADE) across training across each method and formulation in the small dataset setting. Learning with stochastic kinematic priors aids with faster learning. Compared to the baseline (blue), all models employing kinematic priors converge much more quickly. Formulation 4, acceleration and steering, converges most quickly. While DKM also converges very quickly, we point out in Table II that the final performance is much poorer.
@@ V-A Performance on Waymo Motion Prediction Dataset
  • Minimum Average Displacement Error (minADE): The average L2 norm between the ground truth and the closest prediction; $minADE(G)=\min_{i}\frac{1}{T}\sum^{T}_{t=1}\|\hat{s}^{t}_{G}-s^{it}_{G}\|_{2}$.
  • Minimum final displacement error (minFDE): The L2 norm between only the positions at the final timestep, $T$; $minFDE(G)=\min_{i}\|\hat{s}^{T}_{G}-s^{iT}_{G}\|_{2}$.
@@ V-B Performance on a Smaller Dataset Setting
We examine the effects of kinematic priors on a smaller dataset size. This is motivated by the fact driving datasets naturally have an imbalance of scenarios, where many samples are representative of longitudinal straight-line driving or stationary movement, and much less are representative of extreme lateral movements such as U-turns. Thus, large and robust benchmarks like the Waymo, Nuscenes, and Argoverse datasets are necessary for learning good models.
However, large datasets are not always accessible depending on the setting. For example, traffic laws, road design, and natural dynamics vary by region. It would be infeasible to expect the same scale and robustness of data from every scenario in the world, and thus trajectory forecasting will run into settings with less data available.
@@
Compared to the results from Table I, the effects of kinematic priors in learning are more pronounced. Since kinematic priors naturally relate the position at one timestep to the position at the next, we believe that the performance boost can be attributed to this natural relationship. In backpropagation, optimization of one position further into the time horizon directly influences predicted positions at earlier timesteps via the kinematic model. Without the kinematic prior, the relation between timesteps may be implicitly related through neural network parameters. When the model lacks data to form a good model of how an agent moves through space, the kinematic model can help to compensate by modeling simple constraints.
@@ VI Discussion and Conclusion
In this paper, we present a simple and easy-to-implement method for including kinematic priors in probabilistic trajectory forecasting. Kinematic priors can also be implemented for deterministic methods where linear approximations are not necessary. With no additional overhead, kinematic priors not only show improvement in models trained on robust datasets but also in suboptimal settings with small datasets and noisy trajectories, with up to $12\%$ improvement in smaller datasets and $1\%$ less performance degradation in the presence of noise for the full Waymo dataset. For overall performance improvement, we find Formulation 1 in Section IV-B1 with velocity components to be the most beneficial and well-rounded to prediction performance.
When there is large-scale data to learn a good model of how vehicles move, we observe that the effects of kinematic priors are less pronounced. This is demonstrated by the less obvious improvements over the baseline in Table I compared to Table II; model complexity and dataset size will eventually out-scale the effects of the kinematic prior. With enough resources and high-quality data, trajectory forecasting models will learn to “reinvent the steering wheel”, or implicitly learn how vehicles move via the complexity of the neural network. For practical autonomous systems, portability and adaptability are important problems to consider aside into the future. We hope that kinematic priors can be considered in trajectory forecasting and simulation as deployment ventures past well-represented urban scenarios to suboptimal settings, whether it is with noisy sensors, smaller datasets to adapt from, or downscaled models for inference speed and portability.
One limitation of our work is that we primarily explore analysis in one-shot settings; future work focusing on kinematic priors for autoregressive approaches would be interesting in comparison to one-shot models with kinematic priors, especially since autoregressive approaches model predictions conditionally based on previous timesteps.

### v3

@@ Abstract
In trajectory forecasting tasks for traffic, future output trajectories can be computed by advancing the ego vehicle’s state with predicted actions according to a kinematics model. By unrolling predicted trajectories via time integration and models of kinematic dynamics, predicted trajectories should not only be kinematically feasible but also relate uncertainty from one timestep to the next. While current works in probabilistic prediction do incorporate kinematic priors for mean trajectory prediction, variance is often left as a learnable parameter, despite uncertainty in one time step being inextricably tied to uncertainty in the previous time step. In this paper, we show simple and differentiable analytical approximations describing the relationship between variance at one timestep and that at the next with the kinematic bicycle model. In our results, we find that encoding the relationship between variance across timesteps works especially well in unoptimal settings, such as with small or noisy datasets. We observe up to a 50% performance boost in partial dataset settings and up to an 8% performance boost in large-scale learning compared to previous kinematic prediction methods on SOTA trajectory forecasting architectures out-of-the-box, with no fine-tuning.
@@ I Introduction
Figure 1: Motivating example for probabilistic kinematic priors. In real-world traffic, vehicles have a constrained range of behaviors. For instance, it is not possible for a vehicle to move side to side directly, and driving in reverse is highly unlikely. Without kinematic priors, neural networks may search the space of all trajectories, possible or impossible. Without accounting for analytical variances in trajectories as in previous work, the range of possible future trajectories may also be unrealistic.
@@
    A simple and effective method for incorporating analytically-derived kinematic priors into probabilistic models for trajectory forecasting (Section IV), which boosts performance and generalization (Section V) and requires trivial additional overhead computation;
@@ II-B Physics-based Priors for Learning
In our method, we pursue the latter. Since we are not modeling complex systems such as cloth or fluid, simulation of traffic agent states require only a simple, fast, and differentiable update. In addition, since the kinematic models do not describe interactions between agents, the complexity of the necessary model is greatly reduced. In this paper, we hypothesize that modeling the simple kinematics (e.g., how a vehicle moves forward) with equations will allow for greater modeling expressivity on behavior.
@@ IV-B1 Formulation 1: Velocity Components
Figure 4: Average Displacement Error (ADE) across training across each method and formulation in the small dataset setting. Learning with stochastic kinematic priors aids with faster learning. Compared to the baseline (blue), all models employing kinematic priors converge much more quickly. Formulation 4 (brown) converges most quickly. While DKM also converges very quickly, we point out in Table II that the final performance is much poorer.
@@ V-A Performance on Waymo Motion Prediction Dataset
  • Minimum Average Displacement Error (minADE): average L2 norm between the ground truth and the closest prediction; $minADE(G)=\min_{i}\frac{1}{T}\sum^{T}_{t=1}\|\hat{s}^{t}_{G}-s^{it}_{G}\|_{2}$.
  • Minimum final displacement error (minFDE): L2 norm between only the positions at the final timestep, $T$; $minFDE(G)=\min_{i}\|\hat{s}^{T}_{G}-s^{iT}_{G}\|_{2}$.
@@ V-B Performance on a Smaller Dataset Setting
We examine the effects of kinematic priors on a smaller dataset size. This is motivated by the fact driving datasets naturally have an imbalance of scenarios, where many samples are representative of longitudinal straight-line driving or stationary movement, and much less are representative of extreme lateral movements such as U-turns. Thus, large and robust benchmarks like the Waymo, Nuscenes, and Argoverse datasets are necessary for learning good models. However, large datasets are not always accessible depending on the setting. For example, traffic laws, road design, and natural dynamics vary by region. It would be infeasible to expect the same scale and robustness of data from every scenario in the world, and thus trajectory forecasting will run into settings with less data available.
@@
Compared to the results from Table I, the effects of kinematic priors in learning are more pronounced. Since kinematic priors analytically relate the position at one timestep to the position at the next, improvements in metrics may suggest that baseline models utilize a large amount of expressivity to model underlying kinematics. In backpropagation, optimization of one position further into the time horizon directly influences predicted positions at earlier timesteps via the kinematic model. Without the kinematic prior, the relation between timesteps may be implicitly related through neural network parameters. When the model lacks data to form a good model of how an agent moves through space, the kinematic model can help to compensate by modeling simple constraints.
@@ VI Discussion and Conclusion
In this paper, we present a simple method for including kinematic relationships in probabilistic trajectory forecasting. Kinematic priors can also be implemented for deterministic methods where linear approximations are not necessary. With nearly no additional overhead, we not only show improvement in models trained on robust datasets but also in suboptimal settings with small datasets and noisy trajectories, with up to $12\%$ improvement in smaller datasets and $1\%$ less performance degradation in the presence of noise for the full Waymo dataset. For overall performance improvement, we find Formulation 1 with velocity components to be the most beneficial and well-rounded to prediction performance.
When there is large-scale data to learn a good model of how vehicles move, we observe that the effects of kinematic priors are less pronounced. This is demonstrated by the less obvious improvements over the baseline in Table I compared to Table II; model complexity and dataset size will eventually out-scale the effects of the kinematic prior. With enough resources and high-quality data, trajectory forecasting models will learn to “reinvent the steering wheel”, or implicitly learn how vehicles move via the complexity of the neural network.
One limitation is that we primarily explore analysis in one-shot prediction; future work focusing on kinematic priors for autoregressive approaches would be interesting in comparison to one-shot models with kinematic priors, especially since autoregressive approaches model predictions conditionally based on previous timesteps.